OpenStudy (anonymous):

find the derivative of f(x)=x^2 + x -3 using the definition of the derivative

OpenStudy (callisto):

\[f'(x)\]\[=\lim_{h \rightarrow 0}\frac{f(x+h) - f(x)}{h}\]\[=\lim_{h \rightarrow 0}\frac{[(x+h)^2 + (x+h) -3] -(x^2 + x -3)}{h}\]Can you simplify the fraction first?

OpenStudy (anonymous):

I used x^2 + 2(x)(h) + h^2 + x +h -3 -x^2 -x +3 = 2x + x ??

OpenStudy (callisto):

x^2 + 2(x)(h) + h^2 + x +h -3 -x^2 -x +3 <- correct for denominator. But it's not equal to 2x + x... \[x^2 + 2(x)(h) + h^2 + x +h -3 -x^2 -x +3\]Group the like terms together: \[x^2 -x^2 + x -x -3 +3+ h^2+ 2(x)(h)+h\] Now, can you simplify the above expression

OpenStudy (anonymous):

basically factor out the h and H(h+2x+1) divided by the h on denominator and = h + 2x + 1?

OpenStudy (callisto):

Yes.

OpenStudy (callisto):

And you got: \[\lim_{h \rightarrow 0} (h+2x+1)\]Can you evaluate the limit?

OpenStudy (anonymous):

Would be 2x + 1 ??? Because h is going to 0

OpenStudy (callisto):

Indeed it is!

OpenStudy (anonymous):

Holy Cow! It came together finally!! Thank you very much for your help! I would have been stuck!!!

OpenStudy (callisto):

You're welcome :)

OpenStudy (anonymous):

Have a great day!!

OpenStudy (callisto):

You too! :)